1) Extreme sports sold 3,870 monthly subscriptions at $39 each for services provided during February. $22,995 was collected in cash and the rest was cold on account.

What is the credit to service Revenue?

2)The accounts payable began with a balance of $520, the account had debits of $961 and credits of $650.

The ending balance of accounts payable is$____
Answer #1

1) The amount credited to service revenue A/c will be = 3870*39

= $ 150,930

The Amount of 22,995 is also included in service revenue .

The journal entry will be

Cash A/c 22,995

Accounts Receivable A/c 127,935

Service Revenue A/c 150,930

2) The begining balance of Accounts payable will be Credit = 520

(-) debit = (961)

(+) credit = 650   

Ending balance ( 520+650-961)    = $ 209

If you are benefited from the solution then please LIKE , if disliked please specify the reason